# Batch learning equals online learning in Bayesian supervised learning

### v1

@@ 5. Probability measures on ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$
From (5.17), noting that the map $\mathfrak{m}^{n}$ is injective, we conclude that $\tilde{\nu}_{X_{n}}=\mathfrak{m}^{n}(\nu_{X_{n}})$ for some $\nu_{X_{n}}\in{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{X_{n}})$. Finally, using Remark 5.6 (2), we conclude the existence of $\nu\in{\mathcal{P}}({\mathcal{Y}})^{{\mathcal{X}}}$ that satisfies (5.13). This completes the proof of Theorem 5.5. ∎
In what follows we shall consider several illustrations of Theorem 5.5. Let $\delta^{\mathcal{X}}:{\mathcal{Y}}^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ is defined as follows
$\delta^{\mathcal{X}}(x):=\delta(x)\in{\mathcal{P}}({\mathcal{Y}})\text{ for }x\in{\mathcal{X}}.$
Since for any $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ the composition $R^{\mathcal{X}}_{X_{m}}\circ\delta^{\mathcal{X}}:{\mathcal{Y}}^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})^{X_{m}}$ is measurable, the map $\delta^{\mathcal{X}}$ is measurable. Thus we can consider the push-forward map
$P_{*}\delta^{\mathcal{X}}:{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})\to{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big).$
We define a map $\hat{\Pi}_{\mathcal{X}}:{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})\to{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ as follows. For any $x\in{\mathcal{X}}$ we set
$\big(\hat{\Pi}_{\mathcal{X}}(\mu)\big)(x):=(R^{{\mathcal{X}}}_{x})_{*}(\mu)\in{\mathcal{P}}({\mathcal{Y}}).$
Corollary 5.7.
Assume that $\mathcal{Y}$} is a Polish space and ${\mathcal{A}}_{\mathcal{Y}}$ is a countable algebra generating ${\mathcal{B}}({\mathcal{Y}})$. Then we have the following commutative diagram
(5.18) ${{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}\delta^{\mathcal{X}}}$$\scriptstyle{\hat{\Pi}_{\mathcal{X}}}$${{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{ev}_{\mathcal{Y}}^{\mathcal{X}}}$${{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}.}$
Consequently, given a projective system of finite sample spaces endowed with first order probability measures
$\Big\{\Big({\Omega}^{X_{m}}_{(A)},\nu_{(A)}^{X_{m}}\in{\mathcal{P}}\big({\Omega}_{|A|}^{X_{m}}\big)\Big):X_{m}\in\mathrm{P_{fin}}({\mathcal{X}}),A\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})\Big\}$
and induced projection maps
$\Big\{(R_{(A),X_{m}}^{X_{n}})_{*}:{\mathcal{P}}({\Omega}_{(A)}^{X_{n}})\to{\mathcal{P}}({\Omega}_{(A)}^{X_{m}}),$
$P_{*}\pi^{(A),X_{m}}_{(B)}:{\mathcal{P}}\big({\Omega}_{(A)}^{X_{m}}\big)\to{\mathcal{P}}\big({\Omega}_{|B|}^{X_{m}}\big)|\colon$
there exists $\nu\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$ such that
(5.19) $\nu_{(A)}^{X_{m}}=P_{*}\pi_{(A)}^{X_{m}}(R^{\mathcal{X}}_{X_{m}})_{*}\nu\in{\mathcal{P}}\big({\Omega}_{(A)}^{X_{m}}\big)$
if and only if there exists $\mu\in{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ such that for any $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ and any $(A)\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})$ we have
(5.20) $\hat{\Pi}_{X_{m}}\big(\nu^{X_{m}}_{(A)}\big)=P_{*}\pi_{(A)}^{X_{m}}R^{{\mathcal{X}}}_{X_{m}}(\mu)\in{\mathcal{P}}({\Omega}_{(A)})^{X_{m}}.$
Remark 5.8.
For the case that $\mathcal{X}$} consists of a single point, the second part of Corollary 5.7 is due to Orbanz [Orbanz2011, Proposition 4.1].
Proof of Corollary 5.7.
To prove the measurability of $\hat{\Pi}_{\mathcal{X}}$ it suffices to show that for any $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ the composition $R^{\mathcal{X}}_{X_{m}}\circ\hat{\Pi}_{\mathcal{X}}:{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})\to{\mathcal{P}}({\mathcal{Y}})^{X_{m}}$ is measurable. Notice that
$R^{\mathcal{X}}_{X_{m}}\circ\hat{\Pi}_{\mathcal{X}}(\mu)=P_{*}R^{\mathcal{X}}_{X_{m}}(\mu).$
Recall that $P_{*}R^{\mathcal{X}}_{X_{m}}$ is measurable by [Giry82, Theorem 1]. It follows that the composition $R^{\mathcal{X}}_{X_{m}}\circ\hat{\Pi}_{\mathcal{X}}$ is measurable.
The last assertion of Corollary 5.7 follows from the first one and Theorem 5.5. ∎
@@ 6.1. MacEachern’s Dependent Dirichlet Processes revisited
Let us first recall the general definition of a DDP from [BJQ2012, Definition 1]. A DDP is generated by a map (stochastic process) $G:{\Omega}\times{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})$ where $({\Omega},P)$ is a probability space and for each $x\in{\mathcal{X}}$ the map $G(\cdot,x):({\Omega},P)\to{\mathcal{P}}({\mathcal{Y}})$ is measurable, or equivalently, the map
@@
    The dependence structure of $\theta_{i}(\cdot,x)$ across $x\in{\mathcal{X}}$ for each $i$ is determined by a family ${\mathcal{C}}_{\mathcal{X}}^{\theta}:=\{C^{\theta}_{x_{1},\ldots,x_{d}}:[0,1]^{d}\to[0,1]\}$ of copula functions describing finite dimenional CDF of $(\hat{\theta}_{i})_{*}P\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$ where $\hat{V}_{i}:{\Omega}\to{\mathcal{Y}}^{\mathcal{X}}$, $\hat{\theta}_{i}({\omega})(x):=\theta_{i}({\omega},x)$.
@@
Furthermore, Theorem 5.5 provides a consistency condition involving a “mean” measure $\mu\in{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$. For the ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$, this corresponds to the map of base measures $G^{0}_{\mathcal{X}}:{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})$ defined by $x\mapsto G^{0}_{x}$. The ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$ is centered around this collection of measures, as ${\mathbb{E}}[G_{x}]=G^{0}_{x}$. The projective system $\{\nu_{(A)}^{X_{m}},X_{m}\in(\mathrm{P_{fin}}({\mathcal{X}}),\leq)\}$ must satisfy the condition (5.14):
$\widehat{ev}_{{\Omega}_{k}}^{(m)}(\nu_{(A)}^{X_{m}})=P_{*}\pi^{X_{m}}_{(A)}R^{\mathcal{X}}_{X_{m}}(G^{0}_{\mathcal{X}}).$
@@
    The corresponding projective system of first-order measures is given by the base measure map $\mu=G^{0}_{\mathcal{X}}:x\mapsto G^{0}_{x}$.
@@
6.2. Posterior distributions and posteriors predictive distributions of Bayesian supervised learning models $\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}},\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$
@@
where ${\mathcal{D}}{\mathcal{P}}(h(t_{i}))\in{\mathcal{P}}({\mathcal{Y}})$ is the marginal Dirichlet process of ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$ evaluated at $t_{i}$. Note that $\mu^{0}_{T_{m},S_{n},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}\in{\mathcal{P}}({\mathcal{Y}})^{m+n}\subset{\mathcal{P}}({\mathbb{R}}^{k})^{m+n}$ has the marginal probability measures defined by ${\mathcal{D}}{\mathcal{P}}(h(t_{i}))$ and ${\mathcal{D}}{\mathcal{P}}(h(t_{i}))$. Thus to determine the joint measure $\mu^{0}_{T_{m},S_{n},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$ it suffices to compute its copula functions $C_{T_{m},S_{n},({\mathbb{R}}^{\mathcal{X}}_{A})_{*}{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$. Knowing $C_{T_{m},S_{n},(R^{\mathcal{X}}_{A})_{*}{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$, we can apply Theorem 3.4.1 in [DS2016] to compute ${\mathcal{P}}_{T_{m}|S_{n},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$, using the recursive formula in Theorem 4.1.
@@
Let $\mathcal{Y}$} be a measurable space, ${\mathcal{X}}:=\{x_{1},\ldots,x_{n}\}$ a finite set, $S_{n}\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$, and $X_{n}=\Pi_{{\mathcal{X}}}(S_{n})\in{\mathcal{X}}^{n}$, $T_{m}\in{\mathcal{X}}^{m}$. Assume that for any $(A)\in\pi({\mathcal{Y}})$ there exists a Markov kernel ${\mathbf{q}}^{n}_{(A),m}:{\mathcal{Y}}^{n}\to{\mathcal{Y}}^{m}$ which is a regular conditional probability measure of the joint distribution of ${\mathbf{p}}^{(A)}_{T_{m},X_{n}}(P_{*}(\pi_{(A)})_{*}^{\mathcal{X}}\mu)$ such that the following diagram is commutative for any $(A)\leq(B)\in\pi({\mathcal{Y}})$.
${{\Omega}_{(B)}^{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\pi^{(B)}_{(A)})^{m}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(B)}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathbf{q}}_{(B),m}^{n}}$$\scriptstyle{(\pi^{(B)}_{(A)})^{n}}$${{\Omega}_{(A)}^{m}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(A)}^{n}.}$$\scriptstyle{{\mathbf{q}}_{(A),m}^{n}}$
Assume that there exists a map ${\mathbf{q}}^{n}_{m}:{\mathcal{Y}}^{n}\to{\mathcal{Y}}^{m}$ such that for any $(A)\in\pi({\mathcal{Y}})$ the following diagram is commutative
${{\mathcal{Y}}^{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\pi_{(A)})^{m}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\mathcal{Y}}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathbf{q}}^{n}_{m}}$$\scriptstyle{(\pi_{(A)})^{n}}$${{\Omega}_{(A)}^{m}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(A)}^{n}.}$$\scriptstyle{{\mathbf{q}}^{n}_{(A),m}}$
@@
This Theorem is proved in the same way as Theorem 3.4, so we omit its proof.

### v2

@@ 5. Probability measures on ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$
Now we consider the following commutative system
(5.14) ${{\mathcal{P}}({\mathcal{Y}})^{X_{n}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{m}^{n}}$$\scriptstyle{(\pi_{(A)})_{*}^{X_{n}}}$${({\mathcal{P}}({\mathcal{Y}}^{X_{n}}),\tilde{\nu}_{X_{n}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}(\pi_{(A)}^{X_{n}})}$${{\mathcal{P}}({\Omega}_{(A)})^{X_{n}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{m}^{n}}$${{\mathcal{P}}\big({\Omega}_{(A)}^{X_{n}},\mathfrak{m}^{n}(\nu^{X_{n}}_{(A)})\big).}$
Lemma 5.7.
The image $\mathfrak{m}^{n}({\mathcal{P}}({\mathcal{Y}})^{X_{n}})$ is a measurable subset of ${\mathcal{P}}({\mathcal{Y}}^{{\mathcal{X}}_{n}})$.
Proof of Lemma 5.7.
Let $\{{\mathcal{A}}_{\mathcal{Y}}\}$ be a countable algebra generating $\Sigma_{\mathcal{Y}}$. Then
$\mathfrak{m}^{n}({\mathcal{P}}({\mathcal{Y}})^{X_{n}})=\Big\{\mu\in{\mathcal{P}}({\mathcal{Y}}^{X_{n}}):\mu(A_{1}\times\ldots\times A_{n})=$
$\Pi_{i=1}^{n}\mu(\underbrace{{\mathcal{Y}}\times\ldots\times{\mathcal{Y}}}_{(i-1)\,times}\times A_{i}\times\underbrace{{\mathcal{Y}}\times\ldots\times{\mathcal{Y}}}_{(n-i)\,times})\text{ for any }A_{i}\in{\mathcal{A}}_{\mathcal{Y}},i\in\overline{1,n}\Big\}$
is a measurable subset of ${\mathcal{P}}({\mathcal{Y}}^{{\mathcal{X}}_{n}})$. This completes the proof of Lemma 5.7. ∎
Completion of the proof of Theorem 5.5. Since the map $\mathfrak{m}^{n}$ is injective, taking into account of Lemma 5.7, we conclude that
$\tilde{\nu}_{X_{n}}=(\mathfrak{m}^{n})_{*}\nu_{X_{n}}$
for some $\nu_{X_{n}}\in{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{X_{n}})$. Noting that the system $\{({\mathcal{P}}({\mathcal{Y}})^{X_{n}},\nu_{X_{n}}),X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})\}$ of probability spaces satisfies the condition of the Kolmogorov extension theorem, we complete the proof of Theorem 5.5. ∎
@@ 6.1. MacEachern’s Dependent Dirichlet Processes revisited
Let us first recall the general definition of a DDP from [BJQ2012, Definition 1]. A DDP is generated by a map (stochastic process)
$G:{\Omega}\times{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})$
where $({\Omega},P)$ is a probability space and for each $x\in{\mathcal{X}}$ the map $G(\cdot,x):({\Omega},P)\to{\mathcal{P}}({\mathcal{Y}})$ is measurable, or equivalently, the map
@@
    The dependence structure of $\theta_{i}(\cdot,x)$ across $x\in{\mathcal{X}}$ for each $i$ is determined by a family ${\mathcal{C}}_{\mathcal{X}}^{\theta}:=\{C^{\theta}_{x_{1},\ldots,x_{d}}:[0,1]^{d}\to[0,1]\}$ of copula functions describing finite dimenional CDF of $(\hat{\theta}_{i})_{*}P\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$ where $\hat{\theta}_{i}:{\Omega}\to{\mathcal{Y}}^{\mathcal{X}}$, $\hat{\theta}_{i}({\omega})(x):=\theta_{i}({\omega},x)$.
@@
Furthermore, Theorem 5.5 provides a consistency condition involving a “mean” measure $\mu\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$. For the ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$, this corresponds to the map of base measures $G^{0}_{\mathcal{X}}:{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})$ defined by $x\mapsto G^{0}_{x}$, regarded as an element in ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\stackrel{{\scriptstyle\mathfrak{m}^{\mathcal{X}}}}{{\hookrightarrow}}{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$. The ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$ is centered around this collection of measures, as ${\mathbb{E}}_{P}[G_{x}]=G^{0}_{x}$. The projective system $\{\nu_{(A)}^{X_{m}},X_{m}\in(\mathrm{P_{fin}}({\mathcal{X}}),\leq)\}$ must satisfy the condition (5.11):
$\widehat{ev}_{{\Omega}_{k}}^{(m)}(\nu_{(A)}^{X_{m}})=P_{*}\pi^{X_{m}}_{(A)}\circ(r^{\mathcal{X}}_{X_{m}})_{*}(\mathfrak{m}^{\mathcal{X}}G^{0}_{\mathcal{X}}).$
@@
    The corresponding projective system of first-order measures is given by the base measure map $\mu=\mathfrak{m}^{\mathcal{X}}\circ G^{0}_{\mathcal{X}}\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$.
@@
6.2. Posterior predictive distributions of Bayesian supervised learning models $\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}},\operatorname{Id}_{{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}},\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$
@@
where ${\mathcal{D}}{\mathcal{P}}(h(t_{i}))\in{\mathcal{P}}({\mathcal{Y}})$ is the marginal Dirichlet process of ${\mathcal{D}}{\mathcal{D}}{\mathcal{P}}$ evaluated at $t_{i}$. Note that $\mu^{0}_{T_{m},S_{n},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}\in{\mathcal{P}}({\mathcal{Y}})^{m+n}\subset{\mathcal{P}}({\mathbb{R}}^{k})^{m+n}$ has the marginal probability measures defined by ${\mathcal{D}}{\mathcal{P}}(h(t_{i}))$ and ${\mathcal{D}}{\mathcal{P}}(h(x_{j}))$. Thus to determine the joint measure $\mu^{0}_{T_{m},S_{n},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$ it suffices to compute its copula functions $C_{T_{m},S_{n},({\mathbb{R}}^{\mathcal{X}}_{A})_{*}{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$. Knowing $C_{T_{m},S_{n},(R^{\mathcal{X}}_{A})_{*}{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$, we can apply Theorem 3.4.1 in [DS2016] to compute ${\mathcal{P}}_{T_{m}|S_{n},{\mathcal{D}}{\mathcal{D}}{\mathcal{P}}}$, using the recursive formula in Theorem 4.1.
@@
Let $\mathcal{Y}$} be a measurable space, ${\mathcal{X}}:=\{x_{1},\ldots,x_{n}\}$ a finite set, $S_{n}\in({\mathcal{X}}\times{\mathcal{Y}})^{n}$, and $X_{n}=\Pi_{{\mathcal{X}}}(S_{n})\in{\mathcal{X}}^{n}$, $T_{m}\in{\mathcal{X}}^{m}$. Assume that for any $(A)\in\pi({\mathcal{Y}})$ there exists a Markov kernel ${\mathbf{q}}^{n}_{(A),m}:{\Omega}_{(A)}^{n}\to{\mathcal{P}}({\Omega}_{(A)}^{m})$ which is a regular conditional probability measure of the joint distribution of ${\mathbf{p}}^{(A)}_{T_{m},X_{n}}(P_{*}(\pi_{(A)})_{*}^{\mathcal{X}}\mu)$ such that the following diagram is commutative for any $(A)\leq(B)\in\pi({\mathcal{Y}})$.
${{\mathcal{P}}({\Omega}_{(B)}^{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}(\pi^{(B)}_{(A)})^{m}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(B)}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathbf{q}}_{(B),m}^{n}}$$\scriptstyle{(\pi^{(B)}_{(A)})^{n}}$${{\mathcal{P}}({\Omega}_{(A)}^{m})}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(A)}^{n}.}$$\scriptstyle{{\mathbf{q}}_{(A),m}^{n}}$
Assume that there exists a Markov kernel ${\mathbf{q}}^{n}_{m}:{\mathcal{Y}}^{n}\to{\mathcal{P}}({\mathcal{Y}}^{m})$ such that for any $(A)\in\pi({\mathcal{Y}})$ the following diagram is commutative
${{\mathcal{P}}({\mathcal{Y}}^{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}(\pi_{(A)})^{m}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\mathcal{Y}}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathbf{q}}^{n}_{m}}$$\scriptstyle{(\pi_{(A)})^{n}}$${{\mathcal{P}}({\Omega}_{(A)}^{m})}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces{\Omega}_{(A)}^{n}.}$$\scriptstyle{{\mathbf{q}}^{n}_{(A),m}}$
@@
This Theorem is proved in the same way as Theorem 3.4, so we omit its proof. s